P(m)=X^3+mx2-x+12 when P(m) is divided by (x-1) the remainder is r. When P(m) is divided by (x-4) the remainder is 8r. Determine the values of m and r


1
Expert's answer
2021-08-10T14:27:39-0400
"(1)^3+m(1)^2-1+12=r"

"(4)^3+m(4)^2-4+12=8r"


"m=r-12"

"2m=r-9"

"m=3"

"r=15"